If length of a rectangle is 5 times its width and if its area is 180 m2 its perimeter is?
Answer by mananth(16946)

width =x
length = 5x
Area = 5x*x=5x^2
5x^2=180
/5
x^2=36
x=6
width = 6 m
length = 30m